# Do It Yourself: Learning Semantic Correspondence from Pseudo-Labels

### v2

@@ Abstract
Finding correspondences between semantically similar points across images and object instances is one of the everlasting challenges in computer vision. While large pre-trained vision models have recently been demonstrated as effective priors for semantic matching, they still suffer from ambiguities for symmetric objects or repeated object parts. We propose to improve semantic correspondence estimation via 3D-aware pseudo-labeling. Specifically, we train an adapter to refine off-the-shelf features using pseudo-labels obtained via 3D-aware chaining, filtering wrong labels through relaxed cyclic consistency, and 3D spherical prototype mapping constraints. While reducing the need for dataset specific annotations compared to prior work, we set a new state-of-the-art on SPair-71k by over 4% absolute gain and by over 7% against methods with similar supervision requirements. The generality of our proposed approach simplifies extension of training to other data sources, which we demonstrate in our experiments.
@@ 1 Introduction
Recently, foundation model features have demonstrated surprisingly high zero-shot performance for this task [49, 62, 63]. However, these features still inhabit various weaknesses for finding correspondences, such as ambiguities for similar object parts [35] or symmetric objects [63]. In addition, supervised methods [63] indicate that there are more effective strategies to combine foundational features for solving semantic correspondence than simple concatenation [62] or weighted averaging [35]. Recent works [33, 63, 59] train models in a supervised manner using manual keypoint annotations and achieve strong performance when evaluated on the same dataset. However, as the required manual annotations are scarce and difficult to obtain, this strategy is not scalable to larger, more diverse datasets. It, thus, remains an open challenge how to best extract and enhance knowledge encoded in foundation models for finding semantic correspondences without using labor-intensive keypoint supervision.
@@
We address both limitations in our work by learning features that significantly outperform the previous SOTA without the requirement for manual keypoint label definitions: Our work demonstrates that training with self-generated labels, i.e., pseudo-labels, where label quality is improved through weak 3D-aware supervision, is surprisingly effective for improving the semantic correspondence performance, also generalizing to heavy appearance changes (Fig. 1). To achieve this, we train a light-weight adapter supervised by the pseudo-labels, which effectively refines foundation model features, see Fig. 2. Specifically, we pursue a zero-shot approach [62] for generating pseudo-labels for image pairs exhibiting moderate viewpoint variation, a strategy that has been demonstrated to be effective in such contexts but fails for larger viewpoint variations [63]. We compose these labels over multiple image pairs to acquire higher label quality for harder correspondence pairs, i.e., with larger viewpoint variance. We then employ a spherical prototype as a weak geometric prior [35] to reject wrong matches, which addresses the inherent challenge of matchability and reduces feature ambiguity.
@@
  • We show that pseudo-labeling is effective for learning better semantic correspondence features and we demonstrate that the quality of pseudo-labels can be improved through 3D-aware chaining, relaxed cyclic-consistency constraints, and the integration of a weak geometric prior.
  • We show that our strategy can be scaled to larger datasets with 3D annotations, further improving the model’s performance.
@@ 2 Related Work
While a simple nearest-neighbor search in feature space proves to be a powerful zero-shot method for correspondence estimation, prior work also uncovered systematic problems of this approach, e.g., in the disambiguation of symmetric object parts [33, 63, 35, 29, 55, 46]. Therefore, several works proposed to append adapter modules that are fine-tuned using supervision with ground truth correspondences [63, 59]. While reducing the labeling requirements, methods that build a joint atlas for objects in multiple images usually require larger quantities of input images and perform test-time optimization [11, 39].Zhang et al. [63] use keypoint-specific information to disambiguate left-right symmetries at test time. While their strategy is simple, not all symmetries can be solved by image transformations, e.g., flipping, and the required keypoint-specific information is generally not available. Fundel et al. [9] proposed fine-tuning distilled foundational features for geometric matching using 3D ground truth data. This approach is, however, inherently limited to the same object instance, resulting in suboptimal results for cross-instance matching. Mariotti et al. [35] proposed learning a map from features to the surface of a sphere using 3D pose information as a weak supervision signal. While this method successfully resolves most symmetries for simple objects like cars, a spherical prior does not accurately capture objects with more complex topology.
@@
Pseudo-labeling. Lee et al. [27] proposed pseudo-labeling, also referred to as self-training, as a technique for semi-supervised training of neural networks, where a small initial set of labels is propagated to a larger set of unlabeled instances. Various methods have been built on top of this concept, showcasing the generality and scalability of this approach [3, 57, 5, 20, 13, 4], sometimes in connection with knowledge distillation from other pre-trained models. Previous works expand manual keypoint annotations and show improved semantic correspondence performance [21, 28, 18]. In our work, we propose to extract pseudo-labels using off-the-shelf foundational features (at training time), which we use to fine-tune a light-weight adapter for the task of semantic correspondence estimation.
@@ 3 Method
Foundational features show strong semantic awareness for zero-shot matching across instances. However, while the high feature similarity of similar object parts is reasonable from a semantic perspective, it is important to distinguish these elements to find accurate semantic correspondences. The clear performance gap between supervised and unsupervised semantic matching methods [63] shows that a learned combination and refinement of features outperforms a simple concatenation or averaging of multiple foundational features. As we want to supervise a refinement stage without using human annotations, we find motivation in the effectiveness of pseudo-labeling for other tasks [20, 19, 4] and hypothesize that such a strategy can also be conducive to improving features for semantic correspondence.
In this section, we first formalize the semantic correspondence task (Sec. 3.1). We then introduce our method which consists of two stages: First, pseudo-labels are generated via a zero-shot strategy with foundational features and subsequent filtering (Sec. 3.2). Then, these pseudo-labels are used to supervise the training of a light-weight adapter $f_{p}$ (Sec. 3.3). Finally, we detail how our method can be scaled to a larger dataset in Sec. 3.4.
@@ 3.1 Problem Definition
However, in contrast to zero-shot strategies that solely apply pre-trained feature extractors [62], we aim to improve the features such that the refined features are better suited for finding semantic correspondences. Since the features $\mathcal{F}^{\text{DINO}}$ [40] and $\mathcal{F}^{\text{SD}}$ [45] generalize well due to their large training corpus, we learn a light-weight adapter $\mathcal{F}=f_{\text{p}}(\tilde{\mathcal{F}})$ on top of the concatenated features ~$\tilde{\mathcal{F}}=\left[\mathcal{F}^{\text{DINO}},\mathcal{F}^{\text{SD}}\right]$, as proposed by Zhang et al. [63].
@@ 3.2 Generation of High-Quality Pseudo-Labels
$\Delta v$ $[0^{\circ},45^{\circ})$ $(0^{\circ},90^{\circ})$ $(45^{\circ},135^{\circ})$ $(90^{\circ},180^{\circ})$ $(135^{\circ},225^{\circ})$
@@
Table 1: Zero-shot performance (SD+DINO) degrades for larger viewpoint changes. The SPair-71k dataset has $8$ azimuth bins with an equidistant size of $45^{\circ}$. We compute the ratio of correctly corresponded keypoints (PCK@0.1 per-img) for increasing viewpoint changes averaged over all categories.
@@
The zero-shot approach using features from Stable Diffusion [45] and DINOv2 [40] performs well for objects that do not heavily change in their topology and appearance and that are shown from a similar viewpoint. However, the performance clearly degrades for larger viewpoint changes, as studied by Zhang et al. [63]. We study the sensitivity of the semantic correspondence performance for varying viewpoint differences using the azimuth angle annotations of the object-centric semantic correspondence dataset SPair-71k [36] in Tab. 1. We find that the zero-shot performance decreases for larger viewpoint changes, i.e., if the viewpoint’s azimuth angle $\phi$ varies by more than $45^{\circ}$.
@@
To account for this, we propose propagating matches through a $K$-tuple $(I_{1},...,I_{K})$, where each pair $(I_{k},I_{k+1})$ represents an image pair of the same category with a small but non-zero viewpoint variation (Fig. 3, (1)), such that $d_{\text{circ}}(\phi_{k},\phi_{k+1})<90^{\circ}$ with the spherical distance $d_{\text{circ}}(\cdot)$, which is equivalent to one azimuth bin in the SPair-71K dataset. The matches $\mathcal{K}^{k}=\{(p^{k}_{i},p^{k+1}_{i}):p^{k}_{i}\in\mathcal{P}^{k}\subseteq\mathcal{M}^{k},p^{k+1}_{i}\in\mathcal{P}^{k+1}\subseteq\mathcal{M}^{k+1}\}$ are propagated via the recursive application of the NN operator for the subsequent pairs:
@@
We observe that this cyclic consistency constraint rejects many keypoint tuples $(p_{i}^{s},p_{i}^{t})$, as the zero-shot matching approach has difficulties mapping back to the exact location of $p_{i}^{s}$. For this reason, we relax the cyclic consistency constraint, where we do not enforce an exact consistency but allow a small deviation from the source point:
@@
with the threshold $\theta_{\text{th}}$. We select the threshold $\theta_{\text{th}}<0.15\cdot\pi$ that allows disambiguating left and right and potentially repeated object parts, such as the wheels of a car. While this thresholding might also remove correct matches for classes that are not represented well by a rigid object prototype, we observe that the number of matches is still large enough to serve as a sufficiently dense supervision signal. Most importantly, only removing pseudo-labels does not modulate the original zero-shot matches. This is in contrast to regularizing with spherical mapper features, which deteriorates the localization accuracy of the original foundational features [35].
@@ 3.3 Supervised Training with Pseudo-Labels
with a window soft-argmax [63, 23] and Gaussian noise samples $\epsilon$. The dense loss is particularly important for us since it propagates the gradient also to areas in the feature map without labels, i.e., unmatched areas. This learned combination and refinement of foundational features is a decisive improvement over previously proposed simple concatenation [62, 63] or weighted averaging of features similarities of a weak geometric regularizer [35].
[…]
@@
Table 2: Per-category PCK@0.1 scores (per-keypoint) on SPair-71k. Best and second best are highlighted. When pre-training on ImageNet3D [34], results are improved. SphMap avg is slightly higher than in the original paper where they used macro averaging. We trained using their code. *DistillDIFT was evaluated using their trained checkpoints.
@@ 3.4 Scaling to a Larger Dataset
The comparably small size of the SPair-71k dataset with only 18 categories and 1800 images in total motivates scaling to a larger dataset with more diverse classes and more images. The spherical mapper [35] is potentially applicable to larger object-centric datasets with 3D annotations but the simple geometric prior limits the improvement when applied via weighted average, as proposed by Mariotti et al. [35]. The geometric-aware flipping used by the current SOTA methods [9, 63] is not applicable to other datasets since it requires access to the keypoint label definitions, only presented in correspondence datasets.
In contrast, our work only relies on categories, masks, and 3D object poses as weak supervision signal, which are available at a larger scale. Therefore, our method can be scaled to larger datasets, leading to generalizable semantic correspondence matching by training on objects that are not present in the SPair-71k dataset. We demonstrate this by training our whole pipeline on the recently proposed ImageNet-3D [34] dataset, which has around 86k images. We train the spherical mapper similarly to Mariotti et al. [35], where we use masks acquired from SAM [42]. As more accurate 3D pose labels are present in the ImageNet-3D dataset, we do not supervise the viewpoint direction only using the bin of the azimuth angle, as proposed by Mariotti et al. [35] for the SPair-71k dataset. Instead, we extend their formulation by computing the correlation coefficient on the two-sphere $\mathcal{S}^{2}$ with the accurate pose annotation:
@@
with $\psi_{\mathbf{I}}\in\mathcal{S}^{2}$ and $\psi_{\mathbf{I^{\prime}}}\in\mathcal{S}^{2}$ computed from the 3D pose annotations of the objects in the two images. This not only supervises the azimuth viewpoint information (yaw) but also the pitch of an object and, therefore, improves canonicalization. We provide more details in Sec. A.1.
This strategy enables us to scale our proposed method to more classes and images from a significantly larger dataset, eventually enabling better generalization.
@@
Figure 4: Qualitative results for challenging examples. We present four challenging examples of the SPair-71K test dataset where current SOTA models mostly fail. Correct matches that lie within the PCK@${}_{\text{bbox}}@0.1$ radius of the ground-truth label are color-coded in green, while incorrectly matched keypoints are depicted in red.
@@ 4 Experiments
Applying our sampling strategy, we generate 30k image pairs with pseudo-labels from each category in the SPair-71k dataset. For each image pair, we use a maximum of 50 randomly sampled keypoints for supervised training. For the zero-shot generation of pseudo-labels, we follow Zhang et al. [62] and resize the input images to $960^{2}$ and $840^{2}$ to extract the SD and DINOv2 features, resulting in feature maps of resolution $60\times 60$. We choose the rejection radius $r_{\text{max}}$ for the relaxed bicyclic consistency constraint such that a deviation of one feature patch is allowed.
Details on training with pseudo-labels.
Similar to Zhang et al. [63], we use four bottleneck layers with 5M parameters in total for the adapter. The model is trained using the AdamW optimizer with a weight decay rate $0.001$, a learning rate of $5\cdot 10^{-3}$, and the one-cycle scheduler for 200k steps.
@@
Recent works [49, 62, 63, 35] find SPair-71k [36] to be the most challenging and informative benchmark for semantic matching due to its larger size and more challenging matches, containing images from $18$ different categories with $600$ to $900$ image pairs per category. We, additionally, evaluate the methods on the recently proposed AP-10k [61] semantic correspondence dataset [63] without re-training on it.
@@
We compare our method with recent zero-shot and weakly-supervised strategies. DIFT [49], DINOv2 [40], and SD + DINOv2 [62] extract features from pre-trained foundation models and directly apply nearest-neighbor matching in feature space. DistillDIFT (U.S.) [9] is further fine-tuned for correspondence detection on 3D data. ASIC [11] uses DINO features to establish coarse correspondences and refines them at test-time. Spherical mapper (SphMap[35] uses class labels, object masks, and 3D pose information as weak supervision signals during training, similar to us. We note, however, that we do not use any of this information at test time. Telling Left from Right (TLR[63] and DistillDIFT (W.S.) [9] make use of the keypoint label definitions to flip images at test time. We compare to zero-shot approaches and the spherical mapper Mariotti et al. [35] trained on ImageNet-3D with our extended 2-sphere configuration.

### v3

@@ Abstract
Finding correspondences between semantically similar points across images and object instances is one of the everlasting challenges in computer vision. While large pre-trained vision models have recently been demonstrated as effective priors for semantic matching, they still suffer from ambiguities for symmetric objects or repeated object parts. We propose improving semantic correspondence estimation through 3D-aware pseudo-labeling. Specifically, we train an adapter to refine off-the-shelf features using pseudo-labels obtained via 3D-aware chaining, filtering wrong labels through relaxed cyclic consistency, and 3D spherical prototype mapping constraints. While reducing the need for dataset-specific annotations compared to prior work, we establish a new state-of-the-art on SPair-71k, achieving an absolute gain of over 4% and of over 7% compared to methods with similar supervision requirements. The generality of our proposed approach simplifies the extension of training to other data sources, which we demonstrate in our experiments.
@@ 1 Introduction
Recently, foundation model features have demonstrated surprisingly high zero-shot performance for this task [49, 62, 63]. However, these features still have various weaknesses for finding correspondences, such as ambiguities for similar object parts [35] or symmetric objects [63]. In addition, supervised methods [63] indicate that there are more effective strategies to combine foundational features for solving semantic correspondence than simple concatenation [62] or weighted averaging [35]. Recent works [33, 63, 59] train models in a supervised manner using manual keypoint annotations and achieve strong performance when evaluated on the same dataset. However, as the required manual annotations are scarce and difficult to obtain, this strategy is not scalable to larger, more diverse datasets. It thus remains an open challenge to determine the best way to extract and enhance knowledge encoded in foundation models for finding semantic correspondences without relying on labor-intensive keypoint supervision.
@@
We address both limitations in our work by learning features that significantly outperform the previous SOTA without requiring manual keypoint label definitions: Our work demonstrates that training with self-generated labels, i.e., pseudo-labels, where label quality is improved through weak 3D-aware supervision, is surprisingly effective for improving the semantic correspondence performance, also generalizing to heavy appearance changes (Fig. 1). To achieve this, we train a lightweight adapter supervised by the pseudo-labels, which effectively refines the foundation model’s features, as shown in Fig. 2. Specifically, we pursue a zero-shot approach [62] for generating pseudo-labels for image pairs exhibiting moderate viewpoint variation, a strategy that has been demonstrated to be effective in such contexts but fails for larger viewpoint variations [63]. We compose these labels over multiple image pairs to acquire higher label quality for harder correspondence pairs, i.e., with larger viewpoint variance. We then employ a spherical prototype as a weak geometric prior [35] to reject wrong matches, which addresses the inherent challenge of matchability and reduces feature ambiguity.
@@
  • We show that pseudo-labeling is effective for learning better semantic correspondence features, and we demonstrate that the quality of pseudo-labels can be improved through 3D-aware chaining, relaxed cyclic-consistency constraints, and the integration of a weak geometric prior.
  • We show that our strategy is scalable to a larger dataset, further improving the model’s performance.
@@ 2 Related Work
While a simple nearest-neighbor search in feature space proves to be a powerful zero-shot method for correspondence estimation, prior work also uncovered systematic problems of this approach, e.g., in the disambiguation of symmetric object parts [33, 63, 35, 29, 55, 46]. Therefore, several works have proposed appending adapter modules that are fine-tuned using supervision with ground-truth correspondences [63, 59]. While reducing the labeling requirements, methods that build a joint atlas for objects in multiple images usually require larger quantities of input images and perform test-time optimization [11, 39]. Zhang et al. [63] use keypoint-specific information to disambiguate left-right symmetries at test time. While their strategy is simple, not all symmetries can be solved by image transformations, e.g., flipping, and the required keypoint-specific information is generally not available. Fundel et al. [9] proposed fine-tuning distilled foundational features for geometric matching using 3D ground truth data. This approach is, however, inherently limited to the same object instance, resulting in suboptimal results for cross-instance matching. Mariotti et al. [35] proposed learning a map from features to the surface of a sphere using 3D pose information as a weak supervision signal. While this method successfully resolves most symmetries for simple objects, such as cars, a spherical prior does not accurately capture objects with more complex topologies.
@@
Pseudo-labeling. Lee et al. [27] proposed pseudo-labeling, also referred to as self-training, as a technique for semi-supervised training of neural networks, where a small initial set of labels is propagated to a larger set of unlabeled instances. Various methods have been developed based on this concept, demonstrating the generality and scalability of this approach [3, 57, 5, 20, 13, 4], sometimes in conjunction with knowledge distillation from other pre-trained models. Previous works have expanded manual keypoint annotations and demonstrated improved semantic correspondence performance [21, 28, 18]. In our work, we propose to extract pseudo-labels using off-the-shelf foundational features (at training time), which we use to fine-tune a light-weight adapter for the task of semantic correspondence estimation.
@@ 3 Method
Foundational features show strong semantic awareness for zero-shot matching across instances. However, while the high feature similarity of similar object parts is reasonable from a semantic perspective, distinguishing them is crucial for accurate semantic correspondences. The performance gap between supervised and unsupervised semantic matching methods [63] shows that a learned combination and refinement of features outperforms a simple concatenation or averaging of multiple foundational features. As we aim to supervise a refinement stage without relying on human annotations, we draw on the effectiveness of pseudo-labeling for other tasks [20, 19, 4] and hypothesize its potential for improving features for semantic correspondence.
In this section, we first formalize the semantic correspondence task (Sec. 3.1). Then, we introduce our two-stage method: First, pseudo-labels are generated via a zero-shot matching strategy and subsequently filtered (Sec. 3.2). These pseudo-labels are then used to supervise the training of a light-weight adapter $f_{p}$ (Sec. 3.3). Finally, we detail how our method can be scaled to a larger dataset in Sec. 3.4.
@@ 3.1 Problem Definition
[…]
@@ 3.2 Generation of High-Quality Pseudo-Labels
$\Delta v$ $0^{\circ},45^{\circ})$ $(0^{\circ},90^{\circ})$ $(45^{\circ},135^{\circ})$ $(90^{\circ},180^{\circ})$ $(135^{\circ},225^{\circ})$
@@
Table 1: Zero-shot performance (SD+DINO) degrades for larger viewpoint changes. The SPair-71k dataset has 88 azimuth bins with an equidistant size of $45^{\circ}$. We compute the ratio of correctly corresponded keypoints (PCK@0.1 per-img) for increasing viewpoint changes averaged over all categories.
@@
The zero-shot approach, which utilizes features from Stable Diffusion [45] and DINOv2 [40], performs well for objects that do not undergo significant changes in their topology and appearance, and are presented from a similar viewpoint. However, the performance clearly degrades for larger viewpoint changes, as studied by Zhang et al. [63]. We study the sensitivity of the semantic correspondence performance for varying viewpoint differences using the azimuth angle annotations of the object-centric semantic correspondence dataset SPair-71k [36] in Tab. 1. We find that the zero-shot performance decreases for larger viewpoint changes, i.e., if the viewpoint’s azimuth angle $\phi$ varies by more than $45^{\circ}$.
@@
To account for this, we propose propagating matches through a $K$-tuple $(I_{1},...,I_{K})$, where each pair $(I_{k},I_{k+1})$ represents an image pair of the same category with a small but non-zero viewpoint variation (Fig. 3, (1)), such that $d_{\text{circ}}(\phi_{k},\phi_{k+1})<90^{\circ}$ with the spherical distance $d_{\text{circ}}(\cdot)$, which is equivalent to one azimuth bin in the SPair-71K dataset. The matches $\mathcal{K}^{k}=\{(p^{k}_{i},p^{k+1}_{i}):p^{k}_{i}\in\mathcal{P}^{k}\subseteq\mathcal{M}^{k},p^{k+1}_{i}\in\mathcal{P}^{k+1}\subseteq\mathcal{M}^{k+1}\}$ are propagated via recursive application of the NN operator for the subsequent pairs:
@@
We observe that this cyclic consistency constraint rejects many keypoint tuples $(p_{i}^{s},p_{i}^{t})$, as the zero-shot matching approach has difficulties mapping back to the exact location of $p_{i}^{s}$. For this reason, we relax the cyclic consistency constraint, where we do not enforce exact consistency but allow a small deviation from the source point:
@@
with the threshold $\theta_{\text{th}}<0.15\cdot\pi$ that is selected to allow disambiguating left and right and potentially repeated object parts, such as the wheels of a car. While this thresholding may also remove correct matches for classes that are not well-represented by a rigid object prototype, we observe that the number of matches is still sufficiently large to serve as a dense supervision signal. Most importantly, only removing pseudo-labels does not modulate the original zero-shot matches. This is in contrast to regularizing with spherical mapper features, which deteriorates the localization accuracy of the original foundational features [35].
@@ 3.3 Supervised Training with Pseudo-Labels
with a window soft-argmax [63, 23] and Gaussian noise samples $\epsilon$. The dense loss is particularly important for us since it propagates the gradient also to areas in the feature map without labels, i.e., unmatched areas. This learned combination and refinement of foundational features is a decisive improvement over previously proposed simple concatenation [62, 63] or weighted averaging of feature similarities of a weak geometric regularizer [35].
M v [Uncaptioned image] Ω T [Uncaptioned image] [Uncaptioned image] avg
@@
Table 2: Per-category PCK@0.1 scores (per-keypoint) on SPair-71k. Best and second best are highlighted. When pre-training on ImageNet3D [34], results are improved. SphMap avg is slightly higher than in the original paper where they used macro averaging. We trained using their code. *DistillDIFT was evaluated using their trained checkpoints. SphMap and DIY-SC require 3D pose annotations.
@@ 3.4 Scaling to a Larger Dataset
The comparatively small size of the SPair-71k dataset, with only 18 categories and 1,800 images in total, motivates scaling to a larger dataset with more diverse classes and a greater number of images. The spherical mapper [35] is potentially applicable to larger object-centric datasets with 3D annotations, but the simple geometric prior limits the improvement when applied via weighted average, as proposed by Mariotti et al. [35]. The geometric-aware flipping used by the current SOTA methods [9, 63] is not applicable to other datasets since it requires access to the keypoint label definitions, only presented in correspondence datasets.
In contrast, our work relies solely on categories, masks, and 3D object poses as a weak supervision signal, which are available on a larger scale. Therefore, our method can be scaled to larger datasets, leading to generalizable semantic correspondence matching by training on objects that are not present in the SPair-71k dataset. We demonstrate this by training our whole pipeline on the recently proposed ImageNet-3D [34] dataset, which has around 86k images. We train the spherical mapper similarly to Mariotti et al. [35], where we use masks acquired from SAM [42]. As more accurate 3D pose labels are present in the ImageNet-3D dataset, we do not supervise the viewpoint direction only using the bin of the azimuth angle, as proposed by Mariotti et al. [35] for the SPair-71k dataset. Instead, we extend their formulation by computing the correlation coefficient on the two-sphere $\mathcal{S}^{2}$ with the pose annotation:
@@
with $\psi_{\mathbf{I}}\in\mathcal{S}^{2}$ and $\psi_{\mathbf{I^{\prime}}}\in\mathcal{S}^{2}$ computed from the 3D pose annotations of the objects in the two images. This not only supervises the azimuth viewpoint information (yaw) but also the pitch of an object and, therefore, improves canonicalization. We provide more details in the supplementary.
This strategy enables us to scale our proposed method to more classes and images from a significantly larger dataset, ultimately leading to improved generalization.
@@
[…]
@@ 4 Experiments
Using our sampling strategy, we generate 30k pseudo-labeled image pairs per category from SPair-71k. We use a maximum of 50 randomly sampled keypoints for supervised training for each image pair. Following Zhang et al. [62], we resize the input images to $960^{2}$ and $840^{2}$ to extract the SD and DINOv2 features, resulting in feature maps of resolution $60\times 60$. We choose the rejection radius $r_{\text{max}}$ for the relaxed bicyclic consistency constraint such that a deviation of one feature patch is allowed.
Details on adapter training.
Similar to Zhang et al. [63], we use four bottleneck layers with 5M parameters for the adapter. The model is trained using the AdamW optimizer with a weight decay rate $0.001$, a learning rate of $5\cdot 10^{-3}$, and the one-cycle scheduler for 200k steps.
@@
Recent works [49, 62, 63, 35] find SPair-71k [36] to be the most challenging and informative semantic matching benchmark due to its larger size and more challenging matches, containing images from $18$ different categories with $600$ to $900$ image pairs per category. We also evaluate the methods on the recently proposed AP-10k [61] semantic correspondence dataset [63] without training on it.
@@
We compare our method with recent zero-shot and weakly-supervised strategies. DIFT [49], DINOv2 [40], and SD + DINOv2 [62] extract features from foundation models and perform nearest-neighbor matching in feature space. DistillDIFT (U.S.) [9] is further fine-tuned for correspondence detection on 3D data. ASIC [11] uses DINO features to establish coarse correspondences and refines them at test-time. Spherical mapper (SphMap[35] uses class labels, object masks, and 3D pose information as weak supervision signals during training, similar to us. We note, however, that we do not use any of this information at test time. Telling Left from Right (TLR[63] and DistillDIFT (W.S.) [9] make use of the keypoint label definitions to flip images at test time. We compare to zero-shot approaches and the spherical mapper [35] trained on ImageNet-3D with our extended 2-sphere configuration.